OpenStudy (anonymous):

I will be really grateful if someone can explain this to me. A battery connected in series with a resistor R of resistance 5.0 Ω. The electromotive force (e.m.f.) of the battery is 9.0 V and the internal resistance is r. The potential difference (p.d.) across the battery terminals is 6.9 V. Calculate the current in the circuit.

OpenStudy (anonymous):

Now the formula I used was E=V+Vr where V= terminal p.d and Vr=p.d across battery 9=6.9+I(5) I= 0.42A But apparently the answer is incorrect. Why? @experimentX @ganeshie8

OpenStudy (sidsiddhartha):

but it seems correct

OpenStudy (sidsiddhartha):

E-ir=IR

OpenStudy (anonymous):

No, the answer is 1.38 A.

OpenStudy (anonymous):

I know. That's what is confusing me. The formula is right, then why is the answer incorrect?

OpenStudy (sidsiddhartha):

 that's the ckt is'nt it